OpenStudy (anonymous):

How do i do this? Simplify using only positive exponents: (81/64)^ -1/2

OpenStudy (anonymous):

81 ^-1/2 = 1/9 1/64 ^ -1/2 = 1/1/8 = 8 answer is 8/9

OpenStudy (anonymous):

or (81/64)^ -1/2 = (64/81)^ 1/2 = 64^1/2 / 81^1/2 = 8/9

OpenStudy (anonymous):

\[1/\sqrt{(81/64)}\] \[1/(9/8)\] 8/9

jhonyy9 (jhonyy9):

yes 8/9 is right

OpenStudy (anonymous):

hm..is it 5^x = 125?